A weak acid has a ka of 6.2×10^-4 (a)what are the PH and percent ionization of the in 3×10^-3M solution (b) what is the PH and percent ionization of the acid when the solution in (a) is diluted 1000- fold?
1
Expert's answer
2019-10-23T07:01:26-0400

Q97180


Solution: Let HA be a weak acid with ionization constant "\alpha" .


"HA \leftrightarrow H^+ + A^-" ---(reaction at equilibrium)

"c" ---(concentrations at "t=0")

"c(1-\alpha) c\alpha" "c\alpha" ---(concentrations at "t=t" )


Thus, "K_a=[H^+][A^-]\/[HA]=6.2*10^{-4}" ----(given)

"\implies (c\alpha*c\alpha)/c(1-\alpha)=6.2*10^{-4}"

"\therefore c\alpha^2/(1-\alpha)=6.2*10^{-4}" ----(1)


(a)

Given : "c=3*10^{-3}M"


Using (1), we get;

"(3*10^{-3})\alpha^2+(6.2*10^{-4})\alpha-6.2*10^{-4}=0"


Solving the above quadratic equation for "\alpha" , we get;

"\alpha=0.363" (Answer)

(negative root is neglected as percent ionization is always positive)


Thus, percent ionization is "36.3 %" %


Now, "pH =(-log_{10}[H^+])=(-log_{10}(c\alpha))"

"=-log_{10}(3*10^{-3}*0.363)"

"=2.963" (Answer)


pH of the solution at a concentration of "3*10^{-3}M" is "2.963"



(b)

Now; solution is diluted 1000 times.

Thus, the concentration becomes; "c=3*10^{-6}M"


Again, using (1), we get;

"(3*10^{-6})\alpha^2+(6.2*10^{-4})\alpha-6.2*10^{-4}=0"


Solving the above quadratic equation for "\alpha" we get;

"\alpha=0.995" (Answer)

(negative root is neglected as percent ionization is always positive)


Thus, percent ionization is "99.5" %


Again, "pH =(-log_{10}[H^+])=(-log_{10}(c\alpha))"

"=-log_{10}(3*10^{-6}*0.995)"

"=5.525" (Answer)


pH of the solution at a concentration of "3*10^{-6}M" is "5.525"